OpenStudy (anonymous):

I really need help with this! Find the vertex, focus, directrix, and focal width of the parabola. x = 4y^2

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Hello, BG!! The simplest relevant form of the equation of a parabola is 4px=y^2. Compare that to your x = 4y^2. With some relatively straightforward algebra, we can find the value of p. Can you do that? Note: p represents the distance between focus and vertex.

OpenStudy (anonymous):

So if we just swap them, p=1?

OpenStudy (mathmale):

We need to check that. Substituting your p=1 into 4px=y^2, 4x=y^2. Is this the same as your given equation, x=4y^2?

OpenStudy (anonymous):

Nope, wait. y^2=x/4.

OpenStudy (mathmale):

So what is the correct value of p?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I'm not catching on. Not enough sleep.

OpenStudy (mathmale):

:) Compare 4px=y^2 to your x=4y^2. From the first equation, y^2=4px. Substitute this expression for y^2 into x=4y^2. Then solve the resulting equation for p.

OpenStudy (anonymous):

So I plug in 4px? That gets me 16px... Right?

OpenStudy (mathmale):

You're dealing with equations here, so your result should be an equation. Mind trying once more?

OpenStudy (mathmale):

x=4y^2 becomes x=4(4px) after y^2=4px has been substituted.

OpenStudy (mathmale):

solve this equation for p.

OpenStudy (anonymous):

p=x/16x, so 1/16?